Janina is the secretary to the president and she recieves a basic monthly pay ofP 30 000. She is granted a transportation allowance of P500.00 per month. Her deductions are 3.63% for SSS, P 250.00 for PhilHealth, P 4 167.00 + 25% overP 15 833.00 for withholding tax, and 2% for Pag-ibig. Compute her gross earnings and net pay for the month of November.

1
Expert's answer
2021-05-07T09:18:24-0400

Gross earnings= Basic salary +total allowances

=P 30000 +p 500

=P 30500

Net pay =Gross earnings- total deductions

Deductions include SSS=3.63/100*30500= P 1107.15

Phil Health =P 250.00

Tax =P 4167.00+25%over 15833.00

30500-15833 =14667

25/100*14667= P 3666.75

3666.75+4167 = P7833.75

Page ibig 2/100*30500= P610

Total deductions= 1107.15+250+ 7833.75+610 =P 9800.90

Net pay=30500-9800.90= P 20699.10

Net pay=P 20699.10